OpenStudy (anonymous):

find the y-intercept of 3x-5y=6 my options: -6/5 6 3/5

OpenStudy (saifoo.khan):

When y-intercepts, x = 0.

OpenStudy (anonymous):

So -5y=6?

OpenStudy (anonymous):

Divide by -5?

OpenStudy (nurali):

3x-5y=6 y intercept means x=0 3(0)-5y=6 -5y=6 y=-6/5